OpenStudy (anonymous):

find the ordered pair and solve by the substitution method 0.3-0.2y=4 0.2x+0.3y=5/23

OpenStudy (radar):

Check the equations you have entered, does the first one have x in it"? Looks like it might be 0.3x-0.2y=4

OpenStudy (anonymous):

yes sorry it is 0.3x

OpenStudy (anonymous):

so its 0.3x-0.2y=4 0.2x+0.3y=5/23

OpenStudy (anonymous):

OK SO WHAT WOULD THE ORDERED PAIR BE THEN

OpenStudy (anonymous):

BECAUSE IT SAYS I NEED AN ORDERED PAIR WOW MATH CAN BE SO HARD SOMETIMES

OpenStudy (anonymous):

SO WHAT WOULD THE ANSWER BE SWETTIE

OpenStudy (anonymous):

HOW WOULD I WRITE THAT ANSWER

OpenStudy (anonymous):

OK BECAUSE IT SAID THAT THAT WAS NOT THE RIGHT ANSWER

OpenStudy (radar):

I made error in first attempt. When I eliminated decimals by multiplication of powers of 10,the ordered pair were (9.56522, -5.6522)

OpenStudy (radar):

To get you started change the first equation of 0.3x -0.2y = 4 to be 3x-2y=40 by multiplying all terms by 10. In order to substitute we will express x in terms of y. 3x=2y+40, x=2y/3 +40/3 Remember this as we will substitute it in the second equation. Take the second equation and multiply all terms by 10 getting 2x + 3y= 50/23 We have eliminated the decimals and now will make the substitution. 2(2y/3+40/3) + 3y =50/23 4y/3+80/3 +3y = 50/23 4y/3 +3y = 50/23 - 80/3 13y/3=(150-1840)/69=-1690/69=-24.49 y=(3/13)(-24.49)=-5.652 Now that you have the y value, substitute this value for y one of the equations and solve x. good luck with it.